53.8 grams of a complex, diamminebis(oxalato)cobaltate(II), were dissolved in water and its ligand field splitting energy was measured as 55.0 kJ. What is the color of the complex? Ar: Co = 59; N = 14; H = 1; C = 12; O = 16
1
Expert's answer
2020-06-08T15:50:11-0400

Δo=hc/λ

Where

Δo is the splitting energy

h = plaunk constant 6.627×10(-34)

C= speed of light = 3×10(8)

From the formula

λ= 3.6×10(-30) m

So it has no colour